OpenStudy (anonymous):

Find two numbers whose product is -21, and whose sum is -4.

OpenStudy (anonymous):

Well, a negative and a positive is a negative....

OpenStudy (anonymous):

xy=-21,x+y=-4, x+(-21/x)=-4, x^2+4x-21=0, (x-3)(x+7)=0, x=3 or x = -7, solved

OpenStudy (anonymous):

7 and 3 added are not 4 tho

OpenStudy (anonymous):

7 and 3 would be 10 not 4

OpenStudy (anonymous):

i know so what do i do?

OpenStudy (anonymous):

Well, 21 can't go into 4 so you would probably have to use exponents or add/subtract

OpenStudy (anonymous):

do whats the answeR?

OpenStudy (anonymous):

Figure it out. :) If I could I would help you but I'm doing math homework right now and its getting late. Just try to figure out what plus or minus 4 gets you 21 as an answer.